# From Uncertain to Safe: Conformal Adaptation of Diffusion Models for Safe PDE Control

### v1

@@ Abstract
The application of deep learning for partial differential equation (PDE)-constrained control is gaining increasing attention. However, existing methods rarely consider safety requirements crucial in real-world applications. To address this limitation, we propose Safe Diffusion Models for PDE Control (SafeDiffCon), which introduce the uncertainty quantile as model uncertainty quantification to achieve optimal control under safety constraints through both post-training and inference phases. Firstly, our approach post-trains a pre-trained diffusion model to generate control sequences that better satisfy safety constraints while achieving improved control objectives via a reweighted diffusion loss, which incorporates the uncertainty quantile estimated using conformal prediction. Secondly, during inference, the diffusion model dynamically adjusts both its generation process and parameters through iterative guidance and fine-tuning, conditioned on control targets while simultaneously integrating the estimated uncertainty quantile. We evaluate SafeDiffCon on three control tasks: 1D Burgers’ equation, 2D incompressible fluid, and controlled nuclear fusion problem. Results demonstrate that SafeDiffCon is the only method that satisfies all safety constraints, whereas other classical and deep learning baselines fail. Furthermore, while adhering to safety constraints, SafeDiffCon achieves the best control performance.
@@ 1 Introduction
However, existing deep learning-based methods for PDE control generally overlook safety – meaning ensuring control sequences satisfy predefined constraints, mitigating risks, and preventing hazards (Dawson et al., 2022; Liu et al., 2023a). In real-world applications, controlling PDE systems particularly requires addressing safety concerns (Barros & des Santos, 1998; Argomedo et al., 2013). For instance, in fluid dynamics, minor control errors can cause turbulence or structural damage, while in controlled nuclear fusion, violating safety constraints may lead to catastrophic failures. This lack of safety enforcement remains a major bottleneck in applying machine learning to scientific and engineering problems, posing a critical challenge for deployment in high-stakes applications.
@@
Our main contributions are as follows: (1) We introduce safety constraints into deep learning-based control of PDE systems and propose our method SafeDiffCon. (2) To address significant distribution shifts, we quantify the safety uncertainty of diffusion models and design a conformal adaptation process, which incorporates post-training with a reweighted loss and inference-time fine-tuning to promote the output distribution becoming more optimal and safer. (3) We design safe control tasks in the contexts of the 1D Burgers’ equation, 2D incompressible fluid, and the controlled nuclear fusion scenario, and evaluate various methods. Extensive experiments on these datasets demonstrate that SafeDiffCon is the only method to successfully meet all safety constraints while achieving superior control objectives.
@@ 4.1 Uncertainty Quantification of Diffusion Models
The intuition behind the original conformal prediction (Vovk et al., 2005) is to use the trained model on the calibration set to obtain a score set. The score set is then used to endow the model’s prediction with a conformal (confidence) interval. Under the assumption of exchangeability [^2] Exchangeability means that exchanging examples in the calibration set and in inference does not alter their joint distribution., the actual prediction during inference will lie within the conformal interval with a guaranteed probability. The details of conformal prediction are provided in Appendix C.
@@
$\omega(\mathbf{u}_{i},\mathbf{w}_{i})\coloneqq\frac{\tilde{p}(\mathbf{u}_{i},\mathbf{w}_{i})}{p(\mathbf{u}_{i},\mathbf{w}_{i})}$ (7)
@@
$\omega_{\textrm{norm}}(\mathbf{u}_{i},\mathbf{w}_{i})=\frac{\omega(\mathbf{u}_{i},\mathbf{w}_{i})}{\sum_{(\mathbf{u}_{i},\mathbf{w}_{i})\in D_{\text{cal}}}{\omega(\mathbf{u}_{j},\mathbf{w}_{j})}}.$ (8)
@@
$\textrm{CI}_{\theta}(1-\alpha,D_{\textrm{cal}})\coloneqq[s(\mathbf{u}_{\theta}(\mathbf{w}))-Q(1-\alpha;\tilde{\mathcal{S}}),$
$s(\mathbf{u}_{\theta}(\mathbf{w}))+Q(1-\alpha;\tilde{\mathcal{S}})].$ (9)
@@ 4.2 Post-training with Reweighted Loss
The algorithm of post-training is provided in Algorithm 2 in Appendix D.2.
@@ 4.3 Inference-time Fine-tuning
The entire algorithm of inference-time finetuning is provided in Algorithm 3 in Appendix D.3.
@@ 5 Experiment
Figure 2: Visualizations of the 1D Burgers’ equation. The top row shows the original trajectory corresponding to the control target, and the bottom row is the trajectory controlled by SafeConPhy.
@@
with the safety constraint $s_{0}$. If $s(\mathbf{u})>s_{0}$, the state trajectory $\mathbf{u}$ is unsafe, otherwise, it is safe. We compute three unsafe rates to assess the safety levels of different methods’ control results. $\mathcal{R}_{\text{sample}}$ denotes the proportion of unsafe trajectories among total trajectories [^4] If any point in the full trajectory is unsafe, this trajectory is unsafe. So $\mathcal{R}_{\text{sample}}$ is the most stringent metric.; $\mathcal{R}_{\text{time}}$ denotes the proportion of unsafe timesteps among all timesteps; $\mathcal{R}_{\text{point}}$ denotes the proportion of unsafe spatial lattice points in all spatial lattice points across all time steps. More details can be found in Appendix E.1.
@@
Moreover, this control task is particularly challenging due to its specific setup: not only does it require indirect control, which means that control can only be applied to the peripheral region, but the spatial control parameters also reach as many as 1,792. As for safety, among all the training data, 53.1% of the samples are unsafe, meaning their safety score $s$ exceeds the bound $s_{0}=0.1$. The average safety score of the dataset is 0.3215. Other details can be found in Appendix F.1.
@@ 5.3 Tokamak Fusion Reactor
$\mathcal{J}\coloneqq\int_{\Omega\times[0,T]}|\beta_{p}(t,x)-\beta_{p}^{*}(x)|^{2}\mathrm{d}x\mathrm{d}t,$ (18)
@@ 7 Conclusion
Acknowledgements
We gratefully acknowledge the support of Westlake University Research Center for Industries of the Future; Westlake University Center for High-performance Computing. The content is solely the responsibility of the authors and does not necessarily represent the official views of the funding entities.
@@ Appendix A Theory
Appendix B Visualization of Experiment Results
B.1 1D Burgers’ Equation
@@ Appendix B Visualization of Experiment Results
Appendix C Additional Details for Conformal Prediction
@@
Appendix D Details of the Method Implementation
D.1 Conditionally Sample $\mathbf{u}_{\theta}(\mathbf{w})$
@@
D.2 Algorithm of Post-training
@@
D.3 Algorithm of Inference-time Fine-tuning
@@
Appendix E Additional Details for 1D Burgers’ Experiment
E.1 Experiment Setting
@@ E.2 Model
Appendix F Additional Details for 2D Fluid Experiment
F.1 Experiment Setting
@@ F.2 Model
Appendix G Additional Details for Tokamak Fusion Reactor
G.1 Experiment Setting
@@ H.1 CDT
In the 2D incompressible fluid experiment, the model fails to extrapolate safe trajectories with higher returns due to the training data covering a broad range of costs, while the desired cost lies in a narrow range. The augmentation treats all safe trajectories on the Pareto-frontier equally, without emphasizing the region of interest. This complexity, however, could potentially be addressed by SafeConPhy, which is indicated in Table 2. We use the official CDT implementation and follow DT guidelines to sweep desired returns and cost constraints in testing time. In the 1D Burgers experiment, we modify the control objective from the original mean squared error $\mathcal{J}$ between the prediction and target, to an exponential form $\exp\left(-\mathcal{J}\right)$. This new objective is bounded within $[0,1]$, which better aligns with the reward-maximizing setup in reinforcement learning used by CDT. For the 2D incompressible fluid setup, the state, action, and cost prediction heads each consist of 3-layer MLPs with the transformer’s hidden dimension as the inner size.
@@ H.6 TREBI
H.7 PID
Propercentageal Integral Derivative (PID) control (Li et al., 2006) is a versatile and effective method widely employed in numerous control scenarios. For 1D control task, we mainly implement the PID baseline adapted from (Wei et al., 2024). More detailed configurations can be found in Table 14

### v2

@@ Abstract
The application of deep learning for partial differential equation (PDE)-constrained control is gaining increasing attention. However, existing methods rarely consider safety requirements crucial in real-world applications. To address this limitation, we propose Safe Diffusion Models for PDE Control (SafeDiffCon), which introduce the uncertainty quantile as model uncertainty quantification to achieve optimal control under safety constraints through both post-training and inference phases. Firstly, our approach post-trains a pre-trained diffusion model to generate control sequences that better satisfy safety constraints while achieving improved control objectives via a reweighted diffusion loss, which incorporates the uncertainty quantile estimated using conformal prediction. Secondly, during inference, the diffusion model dynamically adjusts both its generation process and parameters through iterative guidance and fine-tuning, conditioned on control targets while simultaneously integrating the estimated uncertainty quantile. We evaluate SafeDiffCon on three control tasks: 1D Burgers’ equation, 2D incompressible fluid, and controlled nuclear fusion problem. Results demonstrate that SafeDiffCon is the only method that satisfies all safety constraints, whereas other classical and deep learning baselines fail. Furthermore, while adhering to safety constraints, SafeDiffCon achieves the best control performance. The code can be found at https://github.com/AI4Science-WestlakeU/safediffcon.
Machine Learning, ICML
@@ 1 Introduction
However, existing deep learning-based methods for PDE control generally overlook safety – meaning ensuring control sequences satisfy predefined constraints, mitigating risks, and preventing hazards (Dawson et al., 2022; Liu et al., 2023a). In real-world applications, controlling PDE systems particularly requires addressing safety concerns (Barros et al., 1998; Argomedo et al., 2013). For instance, in fluid dynamics, minor errors can cause turbulence or structural damage, while in controlled nuclear fusion, violating safety constraints may lead to catastrophic failures. This lack of safety enforcement remains a major bottleneck in applying deep learning to scientific and engineering problems, posing a critical challenge for high-stakes applications.
@@
Our main contributions are as follows: (1) We introduce safety constraints into deep learning-based control of PDE systems and propose our method SafeDiffCon. (2) To address significant distribution shifts, we quantify the safety uncertainty of diffusion models and design a conformal adaptation process, which incorporates post-training with a reweighted loss and inference-time fine-tuning to promote a safer and more optimal output distribution. (3) We design safe control tasks in the contexts of the 1D Burgers’ equation, 2D incompressible fluid, and the controlled nuclear fusion scenario, and evaluate various methods. Extensive experiments on these datasets demonstrate that SafeDiffCon is the only method to successfully meet all safety constraints while achieving superior control objectives.
@@ 4.1 Uncertainty Quantification of Diffusion Models
The intuition behind the original conformal prediction (Vovk et al., 2005) is to use the trained model on the calibration set to obtain a score set. The score set is then used to endow the model’s prediction with a conformal (confidence) interval. Under the assumption of exchangeability [^2] Exchangeability means that exchanging examples in the calibration set and in inference does not alter their joint distribution., the actual prediction during inference will lie within the conformal interval with a guaranteed probability. The details of conformal prediction are provided in Appendix D.
@@
$\omega(\mathbf{u}_{i},\mathbf{w}_{i})\coloneqq\frac{\tilde{p}(\mathbf{u}_{i},\mathbf{w}_{i})}{p(\mathbf{u}_{i},\mathbf{w}_{i})}\vspace{-5pt}$ (7)
@@
$\omega_{\textrm{norm}}(\mathbf{u}_{i},\mathbf{w}_{i})=\frac{\omega(\mathbf{u}_{i},\mathbf{w}_{i})}{\sum_{(\mathbf{u}_{i},\mathbf{w}_{i})\in D_{\text{cal}}}{\omega(\mathbf{u}_{j},\mathbf{w}_{j})}}.\vspace{-5pt}$ (8)
@@
$\vspace{-5pt}\textrm{CI}_{\theta}(1-\alpha,D_{\textrm{cal}})\coloneqq[s(\mathbf{u}_{\theta}(\mathbf{w}))-Q(1-\alpha;\tilde{\mathcal{S}}),$
$s(\mathbf{u}_{\theta}(\mathbf{w}))+Q(1-\alpha;\tilde{\mathcal{S}})].\vspace{-5pt}$ (9)
@@ 4.2 Post-training with Reweighted Loss
The algorithm of post-training is provided in Algorithm 2 in Appendix E.2.
@@ 4.3 Inference-time Fine-tuning
The entire algorithm of inference-time finetuning is provided in Algorithm 3 in Appendix E.3.
@@ 5 Experiment
Figure 2: Visualizations of the 1D Burgers’ equation. The top row shows the original trajectory corresponding to the control target, and the bottom row is the trajectory controlled by SafeDiffCon.
@@
with the safety constraint $s_{0}$. If $s(\mathbf{u})>s_{0}$, the state trajectory $\mathbf{u}$ is unsafe, otherwise, it is safe. We compute three unsafe rates to assess the safety levels of different methods’ control results. $\mathcal{R}_{\text{sample}}$ denotes the proportion of unsafe trajectories among total trajectories [^4] If any point in the full trajectory is unsafe, this trajectory is unsafe. So $\mathcal{R}_{\text{sample}}$ is the most stringent metric.; $\mathcal{R}_{\text{time}}$ denotes the proportion of unsafe timesteps among all timesteps; $\mathcal{R}_{\text{point}}$ denotes the proportion of unsafe spatial lattice points in all spatial lattice points across all time steps. More details can be found in Appendix F.1.
@@
Moreover, this control task is particularly challenging due to its specific setup: not only does it require indirect control, which means that control can only be applied to the peripheral region, but the spatial control parameters also reach as many as 1,792. As for safety, among all the training data, 53.1% of the samples are unsafe, meaning their safety score $s$ exceeds the bound $s_{0}=0.1$. The average safety score of the dataset is 0.3215. Other details can be found in Appendix G.1.
@@ 5.3 Tokamak Fusion Reactor
$\mathcal{J}\coloneqq\int_{\Omega\times[0,T]}(|\beta_{p}(t,x)-\beta_{p}^{*}(x)|^{2}$ (18)
$\quad\quad\quad\quad\quad+|l_{i}(t,x)-l_{i}^{*}(x)|^{2})\mathrm{d}x\mathrm{d}t,$ (19)
@@ 7 Conclusion
Impact Statement
This paper presents work whose goal is to advance the field of Machine Learning. Furthermore, our work makes Machine Learning applications safer in real-world scenarios. There are many potential societal consequences of our work, none of which we feel must be specifically highlighted here.
@@ Appendix A Theory
Appendix B Additional Experiment Results
B.1 Inference Time
We have accelerated the efficiency of inference. On one hand, by introducing post-training, the data distribution is already close to the desired distribution before inference. On the other hand, we have significantly accelerated the diffusion model’s sampling time using DDIM (Song et al., 2021). The Table 5 below presents a comparison of the inference time between our method and the baseline on the Burgers’ equation on A800 with 8 CPUs. It shows that our method achieves a moderate speed and is much faster compared to TREBI, which is also based on the diffusion model.
Table 5: Inference time on the Burgers’ equation on A800 with 8 CPUs.
Methods BC PID SL-Lag MPC-Lag CDT TREBI SafeDiffCon
Inference Time (min) 0.1351 0.1091 0.8842 26.2905 0.0890 13.5525 2.3575
B.2 Ablation Study
In this subsection, we provide results of ablation studies on the 1D Burgers’ equation and the tokamak fusion reactor. We also compare SafeDiffCon and SafeDiffCon without post-training, inference-time fine-tuning, and the uncertainty quantile $Q$. The results in Table 6 and Table 7 still show that the absence of any module affects the model, causing it to fail to meet the safety constraints. Therefore, each module is effective.
Table 6: Results of the ablation study on the Burgers’ equation. We compare SafeDiffCon with SafeDiffCon w/o post-training, inference-time fine-tuning and the uncertainty quantile $Q$.
SafeDiffCon 0.0011 0% 0% 0%
w/o post-training 0.0014 4% 50% 0.01%
w/o fine-tuning 0.0007 40% 20% 3%
w/o $Q$ 0.0006 30% 10% 1%
Table 7: Results of the ablation study on the tokamak fusion reactor. We compare SafeDiffCon with SafeDiffCon w/o post-training, inference-time fine-tuning and the uncertainty quantile $Q$.
SafeDiffCon 0.0094 0% 0%
w/o post-training 0.0153 10% 0.52%
w/o fine-tuning 0.0210 50% 7.74%
w/o $Q$ 0.0269 28% 0.65%
B.3 Analysis on Parameters
We conduct experimental analysis on coverage probability $\alpha$, the split ratio of training/calibration data, and the weight of objective $\gamma$. The Table 8, 9 and 10 show that they do not impact much on performance, indicating that the model is robust.
Table 8: Impact of the coverage probability $\alpha$ on the tokamak fusion reactor.
Split Ratio $\mathcal{J}$downarrow $\mathcal{R}_{\text{sample}}$downarrow $\mathcal{R}_{\text{time}}$downarrow
0.8 0.01017 0% 0%
0.85 0.0091 0% 0%
0.9 0.0094 0% 0%
0.95 0.0095 0% 0%
Table 9: Impact of the split ratio of training/calibration data on the tokamak fusion reactor.
$\alpha$ $\mathcal{J}$downarrow $\mathcal{R}_{\text{sample}}$downarrow $\mathcal{R}_{\text{time}}$downarrow
0.005 0.0124 4% 0.03%
0.01 0.0093 0% 0%
0.02 0.0094 0% 0%
Table 10: Impact of the split ratio of the weight of objective $\gamma$ on the 2D incompressible fluid.
$\gamma$ $\mathcal{J}$downarrow $SVM$downarrow $\mathcal{R}$downarrow
0.01 0.3548 0 0%
0.1 0.4953 0.004 2%
0.3 0.498 0.003 2%
Appendix C Visualization of Experiment Results
C.1 1D Burgers’ Equation
@@ Appendix C Visualization of Experiment Results
Appendix D Additional Details for Conformal Prediction
@@
Appendix E Details of the Method Implementation
E.1 Conditionally Sample $\mathbf{u}_{\theta}(\mathbf{w})$
@@
E.2 Algorithm of Post-training
@@
E.3 Algorithm of Inference-time Fine-tuning
@@
Appendix F Additional Details for 1D Burgers’ Experiment
F.1 Experiment Setting
@@ F.2 Model
Appendix G Additional Details for 2D Fluid Experiment
G.1 Experiment Setting
@@ G.2 Model
Appendix H Additional Details for Tokamak Fusion Reactor
H.1 Experiment Setting
@@ I.1 CDT
In the 2D incompressible fluid experiment, the model fails to extrapolate safe trajectories with higher returns due to the training data covering a broad range of costs, while the desired cost lies in a narrow range. The augmentation treats all safe trajectories on the Pareto-frontier equally, without emphasizing the region of interest. This complexity, however, could potentially be addressed by SafeDiffCon, which is indicated in Table 2. We use the official CDT implementation and follow DT guidelines to sweep desired returns and cost constraints in testing time. In the 1D Burgers experiment, we modify the control objective from the original mean squared error $\mathcal{J}$ between the prediction and target, to an exponential form $\exp\left(-\mathcal{J}\right)$. This new objective is bounded within $[0,1]$, which better aligns with the reward-maximizing setup in reinforcement learning used by CDT. For the 2D incompressible fluid setup, the state, action, and cost prediction heads each consist of 3-layer MLPs with the transformer’s hidden dimension as the inner size.
@@ I.6 TREBI
I.7 PID
Proportional Integral Derivative (PID) control (Li et al., 2006) is a versatile and effective method widely employed in numerous control scenarios. For 1D control task, we mainly implement the PID baseline adapted from (Wei et al., 2024). More detailed configurations can be found in Table 20